Rewrite the equation 2x – 3y = –6 as a function of x
Answer by uma(370)

2x – 3y = –6
adding 3y to both the sides,
2x - 3y + 3y = -6 + 3y
==> 2x = -6 + 3y
Adding 6 to both the sides,
2x + 6 = 3y
==> (2x+6)/3 = y
Thus y = (2x+6)/3 is a function of x.
This is also same as y = 2(x+3)/3
Good Luck!!!